# MedQARo: A Large-Scale Benchmark for Medical Question Answering in Romanian

### v2

@@ 5.2 Preliminary Setup and Experiments
Hyperparameter tuning. We employ consistent hyperparameter settings across all models to ensure a fair comparison among LLM families. All models are fine-tuned using the AdamW Loshchilov and Hutter (2019) optimizer. Due to memory constraints, we set a per-device batch size of one sample with gradient accumulation over $8$ steps, resulting in an effective mini-batch size of $8$ samples. Training is carried out for $5$ epochs, using Brain Floating Point (BFLOAT16) precision to optimize memory usage, while maintaining numerical stability. The learning rate, the dropout rate and the LoRA configuration are established via grid search on the validation set. The range for the initial learning rate is $10^{-3}$ to $10^{-6}$. We employ a cosine learning rate scheduler with $100$ warm-up steps. Dropout is applied with drop rates between 0 and $0.2$, using a step of $0.05$. For the LoRA rank $r$, we look for an optimal value in the set $\{4,6,8,10\}$. For the LoRA scaling factor, we consider values in the set $\{8,16,32\}$. The optimal configuration is based on a learning rate of $2\cdot 10^{-5}$, a dropout rate of $0.05$, a LoRA rank $r$ of $8$, and a LoRA scaling factor $\alpha$ of $16$.
Model Text chunks Validation Test F1 EM BLEU METEOR F1 EM BLEU METEOR Phi-4-mini-instruct First 0.6500 0.6846 0.6366 0.3916 0.6618 0.6935 0.6468 0.3976 Phi-4-mini-instruct All 0.4799 0.5170 0.4106 0.2934 0.4913 0.5288 0.4191 0.2989
@@
Figure 3: Performance per question category (binary, extractive, and reasoning) across four evaluation metrics (F1, EM, BLEU, and METEOR). The values are reported for the top scoring model, namely Phi-4-mini-instruct based on 3,072 tokens, on the official test set. Best viewed in color.
@@
Prompt format. We conduct a preliminary experiment to assess the impact of the prompt structure on QA performance. Specifically, we compare two input formulations for the RoLLaMA2-7B-Instruct model: epicrisis + question + answer (E+Q+A) vs. question + epicrisis + answer (Q+E+A). Both configurations are evaluated with 4,096 input tokens. As shown in Table 4, the Q+E+A format consistently outperforms the E+Q+A format across all evaluation metrics, besides METEOR. Given these results, we decided to run all subsequent experiments using the Q+E+A prompt structure.
5.3 Results
In Table 5, we present the results obtained by the chosen LLMs vs. two baselines on MedQARo. We evaluate all four LLMs under both zero-shot prompting and supervised fine-tuning setups, reporting performance metrics on both validation and test sets.
Effect of prompt length. We test the models with various prompt lengths, taking into account the limitation imposed by each LLM architecture on its input prompt. While the average length of an epicrisis is 7,171 tokens (as per Table 3), the LLMs do not necessarily benefit from using extensive prompts. For example, RoLLaMA2-7B-Instruct, Phi-4-mini-instruct and LLaMA3-OpenBioLLM-8B obtain superior performance levels with 2,048 input tokens than with 4,096 tokens. The only model that obtains better results with 4,096 tokens instead of 2,048 is RoMistral-7B-Instruct, but the performance gains are rather marginal.
Phi-4-mini-instruct is the only model that accepts very long prompts. Although we explore prompts of up to 16,384 tokens for Phi-4-mini-instruct, the model does not benefit from the extended input.
First vs. multiple input chunks. Since epicrises are trimmed from the end to fit the desired prompt lengths, the generally low performance levels when using long prompts can be explained by two factors: (i) most questions can be answered just by looking at the beginning part of an epicrisis, or (ii) taking long contexts into account is a challenging learning task, potentially leading to overfitting. To distinguish between these two cases, we conduct another experiment with one of the best performing models, Phi-4-mini-instruct based on 2,048 input tokens. In addition to the version that simply trims the epicrisis from the end, we evaluate a version that analyzes non-overlapping 2,048-token chunks from the epicrisis. For the latter approach, we employ majority voting over all chunks to obtain the final answer. To break ties, we rely on the average log-probability over the generated tokens and select the answer with the highest value. The corresponding results are reported in Table 6. While the chunk-based approach enables processing of complete epicrises, it results in substantial performance degradation, suggesting that focusing on the first part of the epicrisis leads to optimal results. Still, as shown by our previous experiments reported in Table 5, using a context that is too short, e.g. 1,024 tokens for RoLLaMA2-7B-Instruct or 2,048 tokens for RoMistral-7B-Instruct, might also lead to performance degradation. In summary, trimming the end part of the epicrisis seems to act as a regularization technique, and finding the optimal context length can be assimilated with the process of tuning the corresponding regularization hyperparameter.
@@ 5.3 Results
General remarks. Our experimental results reveal several patterns across all models and configurations. First, we observe that validation and test results remain consistent across all experiments, indicating that our patient-level data splits effectively prevent data leakage. All fine-tuned LLMs exceed the naive baselines by large margins, confirming that models learn meaningful clinical question-answering capabilities. We also find that fine-tuned models significantly outperform their zero-shot counterparts across all architectures, emphasizing that neither prior Romanian language adaptation nor English biomedical pretraining are sufficient for the medical QA task. Among the four LLM families, RoMistral-7B-Instruct obtains the lowest performance levels. In contrast, Phi-4-mini-instruct achieves the highest scores among all evaluated models, despite not being explicitly pretrained on the Romanian language nor on the medical domain. These results confirm our previous observation and similar studies Dorfner et al. (2024), suggesting that prior language or domain adaptation is insufficient.
@@ 5.4 Qualitative Analysis
Analysis of correctly answered questions. In the first example, the LLM was able to recognize the sequence of treatment lines and to distinguish between the first and the second line of therapy, even though the information was not explicitly stated in the epicrisis. In the second example, pulmonary metastases were mentioned under the abbreviation M1PUL. The model recognized that M1PUL denotes the presence of lung metastases, providing the correct answer based on this information. In the third example, the LLM demonstrates the ability to correctly stage breast cancer, which requires high-level medical reasoning. The model correlated tumor size, regional invasion, and the absence of distant metastases, successfully assigning stage IIA, according to the TNM criteria. In the fourth case, the histological type “adenocarcinom” / “adenocarcinoma” was explicitly stated in the histopathology report, which allowed the LLM to correctly answer with “ADK” to the posed question. In the fifth example, the interpretation of the chest CT explicitly stated “invazia tumorală în cadranul supero-extern” / “tumor invasion present in the upper-outer quadrant”. The model was able to accurately spot this information and provide the correct answer. Overall, we observe that the model demonstrates the necessary capabilities to answer both reasoning and extractive questions.
@@
Analysis of wrongly answered questions. In the first example, the annotator considered that the presence of “CHT” in the epicrisis indicates a new line of treatment at disease progression. However, the LLM answered “Nu” / “No”, either because it did not recognize that “CHT” represented a new line of therapy, or because it required an explicit statement confirming the change of treatment line, which is missing from the epicrisis. In the latter case, the model might have considered that “CHT” is part of the initial treatment, instead of a new line of therapy. In the second example, the LLM may have been triggered by the presence of the word “os” / “bone”. The confusion is caused by an investigation reported in the epicrisis, called “scintigrafie osoasă” / “bone scintigraphy”, which is usually recommended for patients with bone metastases. Despite undergoing this investigation, the patient did not have any bone metastases. In the third case, the inflammatory syndrome is not directly mentioned in the epicrisis, although it can be identified by analyzing biological parameters (e.g. elevated ESR, increased CRP, elevated fibrinogen) or secondary diagnoses (e.g. infection, chronic inflammation). The physician recognized these signs and determined the presence of the inflammatory syndrome. Without an explicit mention of “sindrom inflamator” / “inflammatory syndrome”, the LLM was not able to extract the correct answer. In the fourth case, the LLM provided only a partially correct answer. It correctly identified the clinical TNM classification, but did not mention the pathological TNM classification. This is a limitation induced by the prompt length of 3,072 tokens, since the pathological TNM classification is usually specified near the end of the epicrisis, which gets removed due to context trimming. In the fifth example, the patient had lymph node metastases. The physician considered the staging to be stage IV, whereas the LLM interpreted the involvement as loco-regional lymph node invasion and classified the case as stage IIIB. Overall, the wrong or partially incorrect answers point out two kinds of limitations: (i) the reasoning abilities of Phi-4-mini-instruct can sometimes prevent it from integrating high-level medical information, and (ii) the context trimming procedure might sometimes lead to dropping important clues.

### v3

@@ 5.2 Preliminary Setup and Experiments
LoRA adaptation details. All models are fine-tuned using a LoRA strategy applied to the causal language modeling objective. For RoLLaMA2-7B-Instruct, RoMistral-7B-Instruct, and LLaMA3-OpenBioLLM-8B, LoRA adapters are injected into the self-attention projection layers (query, key, value, and output projections), following the default configuration for these architectures. For Phi-4-mini-instruct, due to architectural differences, LoRA is applied to the fused QKV projection layer and the output projection layer, respectively. In all cases, the original model parameters are kept frozen, and only the LoRA adapters are trainable. Feed-forward layers, token embeddings, layer normalization parameters, language modeling heads, and all bias terms remain frozen throughout training. This attention-only adaptation results in updating approximately 0.04-0.10% of the total number of weights, depending on the core architecture.
@@
Hyperparameter tuning. We employ consistent hyperparameter settings across all models to ensure a fair comparison among LLM families. All models are fine-tuned using the AdamW [loshchilov2019decoupled] optimizer. Due to memory constraints, we set a per-device batch size of one sample with gradient accumulation over $8$ steps, resulting in an effective mini-batch size of $8$ samples. Training is carried out for a maximum of $3$ epochs, using Brain Floating Point (BFLOAT16) precision to optimize memory usage, while maintaining numerical stability. The learning rate, the dropout rate and the LoRA configuration are established via grid search on the validation set. The range for the initial learning rate is $10^{-3}$ to $10^{-6}$. We employ a cosine learning rate scheduler with $100$ warm-up steps. Dropout is applied with drop rates between 0 and $0.2$, using a step of $0.05$. For the LoRA rank $r$, we look for an optimal value in the set $\{4,6,8,10\}$. For the LoRA scaling factor, we consider values in the set $\{8,16,32\}$. The optimal configuration is based on a learning rate of $2\cdot 10^{-5}$, a dropout rate of $0.05$, a LoRA rank $r$ of $8$, and a LoRA scaling factor $\alpha$ of $16$.
Table 6: In-domain results of multiple open-source LLMs based on various configurations vs. two trivial baselines, on MedQARo. For each LLM, there is a zero-shot version and several fine-tuned versions, with various prompt lengths. F1, Exact Match (EM), BLEU, and METEOR scores are reported on both validation and test sets. Fine-tuned models consistently outperform zero-shot versions. The best results for each LLM are highlighted in bold green.
Model Fine-tuned #Tokens Validation In-domain test F1 EM BLEU METEOR F1 EM BLEU METEOR Random token selector N/A 1,024 0.0023 0.0006 0.0006 0.0013 0.0019 0.0001 0.0004 0.0011 2,048 0.0021 0.0004 0.0004 0.0013 0.0024 0.0003 0.0005 0.0015 4,096 0.0022 0.0005 0.0006 0.0014 0.0026 0.0003 0.0005 0.0015 Majority answer N/A N/A 0.2091 0.1786 0.1831 0.1153 0.2148 0.1838 0.1884 0.1183 RoLLaMA2-7B 2,048 0.0188 0.0015 0.0024 0.0146 0.0230 0.0025 0.0023 0.0166 1,024 0.5535 0.5431 0.4996 0.3498 0.5646 0.5300 0.4896 0.3451 2,048 0.5832 0.6035 0.5651 0.3487 0.5944 0.6145 0.5780 0.3554 4,096 0.2763 0.2545 0.2508 0.1452 0.3142 0.2890 0.2772 0.1647 RoMistral-7B 4,096 0.0715 0.0430 0.0234 0.0438 0.0605 0.0350 0.0229 0.0403 2,048 0.5761 0.5820 0.5387 0.3425 0.5808 0.5800 0.5331 0.3477 4,096 0.5814 0.5925 0.5487 0.3456 0.5963 0.6045 0.5546 0.3590 Phi-4-mini 3,072 0.0078 0.0000 0.0007 0.0075 0.0048 0.0002 0.0005 0.0043 2,048 0.6500 0.6846 0.6366 0.3916 0.6618 0.6935 0.6468 0.3976 3,072 0.6617 0.6960 0.6472 0.3974 0.6671 0.7011 0.6522 0.4009 4,096 0.4217 0.4145 0.3923 0.2274 0.4220 0.4155 0.3799 0.2303 8,192 0.5050 0.4960 0.4624 0.2911 0.5216 0.5105 0.4653 0.3037 16,384 0.5168 0.5110 0.4765 0.2991 0.5267 0.5130 0.4731 0.3060 32,768 0.5212 0.5182 0.4812 0.3055 0.5158 0.5060 0.4658 0.3067 OpenBioLLM-8B 2,048 0.0140 0.0004 0.0014 0.0156 0.0119 0.0004 0.0015 0.0132 2,048 0.5861 0.6140 0.5720 0.3498 0.5958 0.6215 0.5792 0.3585 4,096 0.5365 0.5855 0.5210 0.3163 0.5566 0.5775 0.5382 0.3348
Prompt format. We conduct a preliminary experiment to assess the impact of the prompt structure on QA performance. Specifically, we compare two input formulations for the RoLLaMA2-7B-Instruct model: epicrisis + question + answer (E+Q+A) vs. question + epicrisis + answer (Q+E+A). Both configurations are evaluated with 4,096 input tokens. As shown in Table 5, the Q+E+A format consistently outperforms the E+Q+A format across all evaluation metrics, besides METEOR. As confirmed by other studies [Barbero-Arxiv-2025], the first tokens tend to receive more attention, which might explain why placing the question at the beginning helps the model focus more on the question that needs to be answered. Given these results, we decided to run all subsequent experiments using the Q+E+A prompt structure.
Table 7: Cross-domain (cross-center and cross-diagnosis) results of multiple open-source LLMs based on various configurations vs. the majority answer baseline, on MedQARo. For each LLM, there is a zero-shot version and several fine-tuned versions, with various prompt lengths. Models are evaluated in terms of F1, EM, BLEU, and METEOR on a single held-out set comprising 231 patients and 3,234 QA pairs. The best results for each LLM are highlighted in bold green.
Model Fine-tuned #Tokens Cross-domain test
F1 EM BLEU METEOR
Majority answer N/A N/A 0.0909 0.0544 0.0544 0.0402
RoLLaMA2-7B 2,048 0.0746 0.0087 0.0138 0.0617
1,024 0.2849 0.2211 0.1979 0.1414
2,048 0.3144 0.2400 0.2325 0.1556
4,096 0.0943 0.0674 0.0427 0.0433
RoMistral-7B 2,048 0.1286 0.0544 0.0469 0.0959
2,048 0.3952 0.3222 0.2815 0.2078
4,096 0.4162 0.3442 0.2864 0.2153
Phi-4-mini 2,048 0.0120 0.000 0.0020 0.0112
2,048 0.3425 0.2934 0.2863 0.1788
3,072 0.3724 0.3142 0.3194 0.1895
4,096 0.1820 0.1555 0.1538 0.0925
8,192 0.3615 0.3077 0.2415 0.1881
16,384 0.3520 0.2981 0.2439 0.1802
32,768 0.3560 0.3012 0.2449 0.1828
OpenBioLLM-8B 2,048 0.0196 0.0006 0.0029 0.0183
2,048 0.3221 0.2764 0.2767 0.1644
4,096 0.2937 0.2492 0.2315 0.1540
5.3 Results
In Table 6, we present the results obtained by the chosen open-source LLMs vs. two baselines on MedQARo. We evaluate all four LLMs under both zero-shot prompting and supervised fine-tuning setups, reporting performance metrics on both validation and test sets.
Table 8: Comparison of best-performing model variants against state-of-the-art baselines on a randomly sampled subset of 100 QA pairs from the in-domain test set. State-of-the-art API-based LLMs are evaluated in the zero-shot setting, while the remaining models are either zero-shot prompted or fine-tuned on MedQARo. The best F1, EM, BLEU, and METEOR scores are highlighted in bold green.
Model Fine-tuned F1 EM BLEU METEOR
GPT-5.2 0.2415 0.1900 0.1339 0.1169
Gemini 3 Flash 0.2041 0.1400 0.0099 0.1039
RoLLaMA2-7B 0.0257 0.0000 0.0009 0.0187
0.5600 0.5800 0.5313 0.3457
RoMistral-7B 0.0892 0.0700 0.0015 0.0568
0.5458 0.5400 0.5011 0.3461
Phi-4-mini 0.0012 0.0000 0.0007 0.0045
0.5950 0.6000 0.5140 0.3440
OpenBioLLM-8B 0.0025 0.0000 0.0003 0.0031
0.5517 0.5900 0.5351 0.3437
Cross-domain evaluation. In Table 7, we report the results obtained on the cross-domain test set, which includes patients from a different medical center, who suffer from other cancer types than those observed during training, thus representing a substantially harder scenario for fine-tuned models. As expected, all fine-tuned models exhibit a noticeable performance drop with respect to the in-domain test set (see Table 6), confirming that cross-center and cross-diagnosis evaluation poses significant challenges. The majority answer baseline achieves low but non-trivial scores, highlighting that naive priors still provide a competitive lower bound in this setting. Zero-shot variants remain largely ineffective, in some cases performing close to the majority answer baseline, which further emphasizes that neither prior Romanian language adaptation nor biomedical pretraining alone is sufficient to handle distribution shifts.
RoMistral-7B-Instruct demonstrates the strongest robustness among all fine-tuned LLMs, achieving the highest F1 and EM scores with both prompt lengths. Phi-4-mini-instruct, while being the top performer in the in-domain setting, shows a more pronounced degradation, indicating that its gains from generic pretraining do not fully translate to unseen diagnostic patterns. Similarly, LLaMA3-OpenBioLLM-8B benefits from fine-tuning, but fails to outperform the Romanian-adapted models, suggesting that English biomedical pretraining does not adequately compensate for cross-lingual and cross-center discrepancies. Among the fine-tuned models, we find that RoMistral-7B attains the highest resilience to both institutional and diagnostic variability. The fact that fine-tuned LLMs remain significantly above their zero-shot counterparts indicates that simple fine-tuning can bring useful question answering capabilities that generalize to different data distributions. Still, fine-tuning alone is not sufficient to achieve full generalization capacity, underlining the opportunity to introduce domain-adaptation strategies.
Effect of prompt length. We test the models with various prompt lengths, taking into account the limitation imposed by each LLM architecture on its input prompt (see Tables 6 and 7). While the average length of an epicrisis is 7,829 tokens (as per Table 3), the LLMs do not necessarily benefit from using extensive prompts. For example, RoLLaMA2-7B-Instruct, Phi-4-mini-instruct and LLaMA3-OpenBioLLM-8B obtain superior performance levels with 2,048 input tokens than with 4,096 tokens. The only model that obtains better results with 4,096 tokens instead of 2,048 is RoMistral-7B-Instruct, but the performance gains on the in-domain test are rather marginal.
Phi-4-mini-instruct is the only model that accepts very long prompts. Although we explore prompts of up to 32,768 tokens for Phi-4-mini-instruct, the model does not benefit from the extended input.
Open-source vs. API-based LLMs. To compare the open-source LLMs with state-of-the-art API-based models, we randomly select 100 QA pairs from the in-domain test set. The limited test set size is determined by our resource constraints, while the use of zero-shot prompting for the API-based models is due to the access restrictions imposed by OpenAI and Google. We present comparative results in Table 8. For a fair comparison, the results of each model are based on an optimal number of input tokens. GPT-5.2 and Gemini 3 Flash substantially outperform the lighter open-source models when both API-based and open-source LLMs are evaluated in the zero-shot setting. This highlights the effectiveness of large-scale general pretraining, even without task-specific supervision. Despite the strong performance of both GPT-5.2 and Gemini 3 Flash, all fine-tuned models consistently outperform the API-based LLMs across all evaluation metrics, confirming the benefits of task-specific and domain-specific adaptation. This comparison reinforces the importance of fine-tuning to achieve high performance on specialized clinical QA tasks in low-resource languages.
@@ 5.3 Results
Model Text chunks Validation In-domain test F1 EM BLEU METEOR F1 EM BLEU METEOR Phi-4-mini First 0.6500 0.6846 0.6366 0.3916 0.6618 0.6935 0.6468 0.3976 All 0.4799 0.5170 0.4106 0.2934 0.4913 0.5288 0.4191 0.2989
First vs. multiple input chunks. Since epicrises are trimmed from the end to fit the desired prompt lengths, the generally low performance levels when using long prompts can be explained by two factors: (i) most questions can be answered just by looking at the beginning part of an epicrisis, or (ii) taking long contexts into account is a challenging learning task, potentially leading to overfitting. To distinguish between these two cases, we conduct another experiment with one of the best performing models, Phi-4-mini-instruct based on 2,048 input tokens. In addition to the version that simply trims the epicrisis from the end, we evaluate a version that analyzes all non-overlapping 2,048-token chunks from the epicrisis. For the latter approach, we employ majority voting over all chunks to obtain the final answer. To break ties, we rely on the average log-probability over the generated tokens and select the answer with the highest value. The corresponding results are reported in Table 9. While the chunk-based approach enables processing of complete epicrises, it results in substantial performance degradation, suggesting that focusing on the first part of the epicrisis leads to optimal results. Still, as shown by our previous experiments reported in Table 6, using a context that is too short, e.g. 1,024 tokens for RoLLaMA2-7B-Instruct or 2,048 tokens for RoMistral-7B-Instruct, might also lead to performance degradation. In summary, trimming the end part of the epicrisis seems to act as a regularization technique, and finding the optimal context length can be assimilated with the process of tuning the corresponding regularization hyperparameter.
Figure 5: Performance per question category (binary, extractive, and reasoning) across four evaluation metrics (F1, EM, BLEU, and METEOR). The values are reported for the top scoring model, namely Phi-4-mini-instruct based on 3,072 tokens, on the in-domain test set. Best viewed in color.
@@
General remarks. Our experimental results reveal several patterns across all models and configurations. First, we observe that in-domain validation and test results remain consistent across all experiments, indicating that our patient-level data splits effectively prevent data leakage. All fine-tuned LLMs exceed the naive baselines by large margins, confirming that models learn meaningful clinical question-answering capabilities. We also find that fine-tuned models significantly outperform their zero-shot counterparts across all architectures, emphasizing that neither prior Romanian language adaptation nor English biomedical pretraining is sufficient for the medical QA task. Among the four open-source LLM families, RoMistral-7B-Instruct obtains the lowest in-domain performance levels. In contrast, Phi-4-mini-instruct achieves the highest scores among all evaluated models on the in-domain test set, despite not being explicitly pretrained on the Romanian language nor on the medical domain. These results confirm our previous observation and similar studies [dorfner2024biomedical], suggesting that prior language or domain adaptation is insufficient. The cross-domain evaluation indicates that fine-tuning is also beneficial for patients enrolled in different medical centers, who suffer from other cancer types. However, fine-tuned LLMs exhibit a visible performance degradation in the cross-center and cross-diagnosis setup, suggesting that fine-tuning alone is not enough to recover the domain gap. Regarding the comparison between the zero-shot API-based LLMs and fine-tuned open-source LLMs, we find that GPT-5.2 and Gemini 3 Flash obtain reasonable performance levels, but they are clearly outperformed by fine-tuned LLMs. This finding emphasizes that fine-tuning open-source LLMs is preferred over zero-shot prompting larger LLMs gated by paywalls.
@@ 5.4 Qualitative Analysis
Analysis of correctly answered questions. In the first example, the LLM was able to recognize the sequence of treatment lines and to distinguish between the first line and the second line of therapy, even though the information was not explicitly stated in the epicrisis. In the second example, pulmonary metastases were mentioned under the abbreviation M1PUL. The model recognized that M1PUL denotes the presence of lung metastases, providing the correct answer based on this information. In the third example, the LLM demonstrates the ability to correctly stage breast cancer, which requires high-level medical reasoning. The model correlated tumor size, regional invasion, and the absence of distant metastases, successfully assigning stage IIA, according to the TNM criteria. In the fourth case, the histological type “adenocarcinom” / “adenocarcinoma” was explicitly stated in the histopathology report, which allowed the LLM to correctly answer with “ADK” to the posed question. In the fifth example, the interpretation of the chest CT explicitly stated “invazia tumorală în cadranul supero-extern” / “tumor invasion present in the upper-outer quadrant”. The model was able to accurately spot this information and provide the correct answer. Overall, we observe that the model demonstrates the necessary capabilities to answer both reasoning and extractive questions.
@@
Analysis of wrongly answered questions. In the first example, the annotator considered that the presence of “CHT” in the epicrisis indicates a new line of treatment at disease progression. However, the LLM answered “Nu” / “No”, either because it did not recognize that “CHT” represented a new line of therapy, or because it required an explicit statement confirming the change of treatment line, which is missing from the epicrisis. In the latter case, the model might have considered that “CHT” is part of the initial treatment, instead of a new line of therapy. Note that this is not exactly a binary question, since the correct answer for this kind of question is a list of treatment lines. When the list is empty, the correct answer is “Nu” / “No”. In the second example, the LLM may have been triggered by the presence of the word “os” / “bone”. The confusion is caused by an investigation reported in the epicrisis, called “scintigrafie osoasă” / “bone scintigraphy”, which is usually recommended for patients with bone metastases. Despite undergoing this investigation, the patient did not have any bone metastases. In the third case, the inflammatory syndrome is not directly mentioned in the epicrisis, although it can be identified by analyzing biological parameters (e.g. elevated ESR, increased CRP, elevated fibrinogen) or secondary diagnoses (e.g. infection, chronic inflammation). The physician recognized these signs and determined the presence of the inflammatory syndrome. Without an explicit mention of “sindrom inflamator” / “inflammatory syndrome”, the LLM was not able to extract the correct answer. In the fourth case, the LLM provided only a partially correct answer. It correctly identified the clinical TNM classification, but did not mention the pathological TNM classification. This is a limitation induced by the prompt length of 3,072 tokens, since the pathological TNM classification is usually specified near the end of the epicrisis, which gets removed due to context trimming. In the fifth example, the patient had lymph node metastases. The physician considered the staging to be stage IV, whereas the LLM interpreted the involvement as loco-regional lymph node invasion and classified the case as stage IIIB. Overall, the wrong or partially incorrect answers point out two kinds of limitations: (i) the reasoning abilities of Phi-4-mini-instruct can sometimes hinder the integration of high-level medical information, and (ii) the context trimming procedure might sometimes lead to dropping important clues.